AN INVESTIGATION OF EXCITED-STATE STRUCTURE AND DYNAMICS IN ACETYLACETONE THROUGH USE OF RESONANCE RAMAN SPECTROSCOPY

\underline{S. A. BROADBENT}, S. M. WILSON, and P. H. VACCARO, Department of Chemistry, Yale University, P.O. Box 208107, New Haven, CT 06520; B. R. JOHNSON, Department of Chemistry, Rice University, P.O. Box 1892, Houston, TX 77251.

Acetylacetone (AA), one of the simplest β-diketones, exhibits a strong intramolecular hydrogen bond that stabilizes the enol tautomer of the isolated (gas-phase) species and mediates the attendant proton- transfer process. Both experiment^{a} and theory^{b} have demonstrated conclusively that the X̃^{1}A_{1} ground electronic state exhibits an asymmetrical equilibrium geometry with a potential barrier of finite height separating two equivalent conformers of C_{s} symmetry. In contrast, ab initio calculations^{c} have suggested that the electronically excited B̃^{1}B_{2} (π^{∗}π) manifold supports a symmetric (C_{2v}) minimum energy configuration which has the shuttling hydron located midway between the oxygen atom centers. This assertion, with its prediction of a low-barrier hydrogen-bonding motif, has been investigated experimentally by means of Resonance Raman Spectroscopy. Excitation at 266 nm, essentially coincident with the peak of the π^{∗} ← π transition, results in Raman profiles dominated by intense spectral features that stem from vibrational modes involving substantial distortion of the chelate ring, including marked displacement of the O⋯O distance. Of special note is the 1620−2800 cm^{−1} region, which is not expected to contain any fundamental transitions, yet exhibits rich structure that has been assigned to overtone and combination bands. All of these data are consistent with a large change in molecular geometry upon electronic excitation. Resonance Raman spectra of deuterated derivatives and structural analogues of AA afford an additional means for unraveling the observed excited-state behavior. Ongoing extensions of these studies will be discussed, as well as efforts toward theoretical analysis based on the time-dependent formalism for Raman scattering.

aR. Boese, M. Y. Antipin, D. Blaser, and K. A. Lyssenko, J. Phys. Chem. B \underline{102}, 1584 (1998).

bV. V. Sliznev, S. B. Lapshina, and G. V. Girichev, J. Struct. Chem. \underline{43}, 47 (2002).

cH. P. Upadhyaya, A. Kumar, and P. D. Naik, J. Chem. Phys. \underline{110}, 11850 (2003).